Quantitative Literacy. Percent Increase or Decrease

Find or make up a situation in which there is an increase or decrease with one piece of your information missing, show how to find the missing part, and show how to check your answer.

example:
According to an article in the newspaper, the state gas tax will be reduced by 1.4 cents next year. If the price per gallon is $3.349, what percent decrease would that be? Solution: $3.349 minus $0.014, equals $3.335. Using the formula (new minus old) / old, we have (3.335 - 3.349)/3.349 = 0.418% decrease. Check: $3.349 + 3.349(-0.00418) = $3.335, the new price.
Answer #1

Consider the following example.

1.Stock in company ABC decreased from $65 a share to $52 a share. What is the percent decrease in stock price.

Percent decrease = (Old Price - New Price)/Old Price * 100% = (65-52)/65 * 100%

                          = 13/65 * 100% = 0.2*100%= 20%

Check:

new price of stock = $65+$65(-0.2) = $65-$13 = $52

2.Jay works in a clothing store for $15 per hour. If his pay is increased to $20 per hour. what is his percent increase in pay?

Percent increase = (New Price - Old Price)/Old Price * 100% = (20-15)/15 * 100%

                          = 5/15 * 100% = 0.33333*100%= 33.33%

Check:

new salary = $15+$15(0.33333) = $15+$4.9999 = $15+$5 = $20